### v1

@@ 7.4 Domain-specific Challenges
IoT and smart systems represent another domain where MI attacks can exploit computationally constrained edge devices. These systems, such as smart home devices or industrial IoT platforms, often lack the computational capacity for traditional, resource-intensive defenses. Lightweight encryption or pruning methods can help mitigate these vulnerabilities, but they must be adapted to the unique resource and latency requirements of edge computing [3].
Future Research: To tackle these challenges, domain-specific defense schemes must be developed. These schemes should involve privacy-preserving techniques supported by an understanding of varying needs (e.g., different data types or resource-limited devices) in each field, ensuring both robustness against MI attacks and usability in real-world scenarios. Focusing on tailored approaches will make future studies better align the goals of privacy, utility, and practicality in diverse application domains.
@@ 8 Conclusion
This survey has provided a comprehensive examination of MI attacks and corresponding defense mechanisms, offering a structured taxonomy of MI attacks stemming from diverse techniques and an in-depth review of their applications across numerous domains, including biometric recognition, healthcare, and finance. After detailed analysis and discussion about various MI defense mechanisms, we highlight the urgent need for robust defenses against MI attacks to secure deep learning systems. Our analysis underscores the vital importance of privacy preservation as a cornerstone for ensuring safe, reliable, and trustworthy use of deep/machine learning, particularly in sensitive applications where data confidentiality is paramount.
This survey presents the challenges of MI-related studies and suggests future research directions, which should prioritize the development of more robust, scalable, and generalizable defense mechanisms that balance privacy preservation with model utility. Future research must also focus on defining realistic threat models, enhancing the explainability of defense techniques, and addressing domain-specific challenges. By fostering collaboration across academia and industry, we can pave the way for developing secure deep learning systems that are protected against MI attacks.

### v2

@@ 7 Challenges and Future Research Directions
While substantial progress has been made in understanding and mitigating MI attacks, many challenges remain. This section discusses some of these challenges and corresponding future research directions.
@@ 7.4 Domain-specific Challenges
IoT and smart systems represent another domain where MI attacks can exploit computationally constrained edge devices. These systems, such as smart home devices or industrial IoT platforms, often lack the computational capacity for traditional, resource-intensive defenses. Lightweight encryption or pruning methods can help mitigate these vulnerabilities, but they must be adapted to the unique resource and latency requirements of edge computing [4]. Given these constraints, conventional encryption techniques may be impractical, increasing the susceptibility of these devices to inference attacks. To address this, lightweight encryption schemes and optimized security protocols are essential, as they can enhance data protection while ensuring efficient transmission and real-time responsiveness [2].
Future Research: To tackle these challenges, domain-specific defense schemes must be developed. These schemes should involve privacy-preserving techniques supported by an understanding of varying needs (e.g., different data types or resource-limited devices) in each field, ensuring both robustness against MI attacks and usability in real-world scenarios. Focusing on tailored approaches will make future studies better align the goals of privacy, utility, and practicality in diverse application domains. In the case of IoT and edge computing, further exploration of lightweight cryptographic methods, secure data transmission protocols, and scalable privacy-preserving techniques is necessary to minimize vulnerabilities without overburdening resource-limited devices [121], especially when attaching AI accelerators to these devices are impractical. Additionally, interdisciplinary efforts that integrate security, privacy, and real-world usability will be critical in ensuring that MI defenses align with the practical needs of different applications.
@@ 8 Conclusion
This survey has provided a comprehensive overview of MI attacks and corresponding defense mechanisms, offering a structured taxonomy of MI attacks based on diverse techniques and an in-depth review of their applications across key domains such as biometric recognition, healthcare, and finance. Through detailed analysis and discussion, we emphasize the increasing importance of developing robust defenses to protect sensitive data in deep learning systems. While summarizing existing defense strategies, we recognize that many current approaches face limitations in terms of scalability, generalizability, and maintaining model utility. Furthermore, most defenses are evaluated under simplified threat models that may not fully capture real-world adversarial capabilities. These limitations highlight the need for continued research in designing more practical and adaptable solutions.
We also identify several promising directions for future work. These include defining realistic threat models, enhancing the explainability of defense techniques, and addressing domain-specific challenges. By tackling these issues, we can pave the way toward developing more secure deep learning systems that are resilient to MI attacks.
[…]
Acknowledgments This research was supported in part by an Australia Research Council (ARC) Discovery Project Grant: DP230102828.